Project List: June 2003

Project Title  Description  Innovations  Completion Time  Most Recent Work

Map of the Web Hypercard program to graphically represent the WWW as a street atlas clever logic for street layout; new way to browse web Convert to C++: 60 hrs; Improve: 60 hrs

Unknown

Binary Op-Code System Command descriptor method that requires only 2 commands; possibly Turing complete Possibly smallest Op-Code system ever; Speed as new MISC processor Prototype Processor: 10 hrs; Create Compiler: 60 hrs

ClockMaker Educational mechanical system creation and modelling for clocks and similar devices Efficient collisiondetection and timeupdating system; sophisticated physicsbased coding structure To simplified beta: 40 hrs; Packaged version: 200 hrs

Evolutionary A.I.

Simple distributed background A.I. that uses G.A. system to predict data from web True A.I. through selfprediction; Combination with web-crawling data-collection Web-Based Packaged Version: 40 hrs

Combined A.I.

Combined Concept-Network, systems, codelet, and G.A. approach to general prediction OO Coding Innovations; Extension of Hofstadter to general A.I.

Destributed version: 40 hrs; Needed Human-Mind Subsystems: 60 hrs

Nathe Programming Language High-level, continuations-based, non-deterministic parallel programming language with programmatic data Bindings as programs; Pattern-matching with non-determinism; Compilation as evaluation; Base types behaviours as infinite definitions Flush-out Language: 25 hrs; Write Compiler: 80 hrs; Write White-Paper: 60 hrs Written Human Language Idea-based, writingmotivated communication structured down to sub-letter level ; Automaticallygenerated dictionary Self-similar sub-letter to sentence structuring; Ability to create arbitrary new words; Speed and beauty in writing; Easy-to-understand writing and speaking system Flush-out Language: 25 hrs; Write Booklet: 100 hrs

Mural-Maker Robot Automated graphical image wall painter with algorithms for depixelizing and pathoptimization Image-to-woodcut decomposition; Automated painting Rebuild Robot: 20 hrs; Improve Program: 40 hrs

Fractal-Planet Project Programs to create and display fractal and evolved planets and EAPS systems Combination of equations and tricks; Multiple-system integrator Packaged System 1.0: 100 hrs

Desktop Sprites Varied, interacting mini-A.I.’s on the desktop, with progressively more sophisticated behaviours Web-based, automated sprite creator; Sprite interaction and evoluation Packaged System 1.0: 80 hrs

Java-Based Organizer Expandable calendar, budget, and eventlogistics tracker, with class and instance hierarchies Java-code based extendability; Integration of organizational systems; Strong structure for tracking organizing objects Base System Completion: 40 hrs; Packaged System 1.0: 80 hrs

Darwin- Compressor Reversible minimal instruction-set applied to evolving compression algorithms; Pentium optimization Reversible, 1 byte opcodes; Evolutionary Optimal Compression Packaged System 1.0: 60 hrs Improved Slide-Rule Any-available function applied or inversed with solid and transparent disconnected siderule Allow any arrangement with disconnect and transparency; Paperthin calculation Packaged System: 40 hrs

Streamboard Keyboard generation based on optimizing “gliding” movement with shortest expected time Efficient typing; Simulated annealing optimization Packaged System: 60 hrs

Role-Playing System More realistic, flexible, and systematic roleplaying system based on 3rd ed. Dungeons and Dragons Less arbitrary than D&D; Skillbased classes; Mental/Physical energy/level system; Less segmented combat Write-up changes for v1.5: 20 hrs; Improve Document: 40 hrs

Web-crawling Background Graphics Ever-changing background graphics with preference-affected web-crawling Changing Background Packaged System: 40 hrs

Automated Chatter Eliza-like chat manager to handle away users with answers to frequent quieries and discussion Tie Eliza-chatter into organization manager Packaged System: 80 hrs

Dual Alarm Clocks One clock with two alarms and other features in two, each password protected, for couples.

Password protected alarms; Doubled alarmclock features Prototyping: 12 hrs

Additional Projects:

1. Economy/Society-structures Simulator Social structure prediction

2. Book of Kings Mental archetype behaviours as people behaviours

3. Preference-based Music Player Multi-tiered Probability-based system with learning of preference-interactions

4. Web-Distributed Full-Featured Interactive Virtual World Efficient message propogation, user-based world creation, many interaction levels

5. Young Person’s Illustrated Primer Story creation from archetypes, plot-elements, and current user needs

6. Rhythm-based A.I.

Trend-finding with Fourier analysis and stimulus-response

7. Evaluation-based A.I.

Learning algorithms in recursive evaluation

8. Miniature Building Blocks Cooperative using of abilities; combine bit of manipulation, memory, logic